OpenStudy (anonymous):

f(x)= ln x(x^2+1)^2/ln(x^3+1)^1/2 differentiate this

OpenStudy (anonymous):

first expand the terms: so u'll get, \[f(x) = /\frac{\ln(x ^{3}+2x ^{2}+x} {\ln \sqrt{x ^{3}+1}}\]

OpenStudy (anonymous):

now use this rule: (u/v)' = (u'v - v'u)/v^2 it'll be a little complicated...but u'll be fine if u do it step-by-step. and i'm here to help.

OpenStudy (anonymous):

thankx :)

OpenStudy (anonymous):

is it ok if u apply both product rule and the quotient rule to solve this ??

OpenStudy (anonymous):

how do u want to apply the product rule?

OpenStudy (anonymous):

product rule for lnx(x^2+1)^2 part and then apply quotient rule to the whole thing

OpenStudy (anonymous):

u can do that...but it'll just complicate things. u can either expand the x(x^2 +1)^2 and then just apply quotient rule...or u can simplify things this way: ln x(x^2 +1)^2 =lnx + ln(x^2+1)^2 =lnx + 2ln(x^2+1) and now just distribute the denominator to both these terms and differentiate.

OpenStudy (anonymous):

i see... thank you for the help (Y)

OpenStudy (anonymous):

sure! happy to help! see if u're getting the right answer.

OpenStudy (anonymous):

oh btw...my first expansion was wrong! i just saw that its x^2...not just x. so the numerator should be: ln(x^5 + 2x^3 +x)